Show or prove that; x ^ (m/n) = n * sqrt(x ^ m) = (root(n, x)) ^ m

1
Expert's answer
2022-03-20T17:17:46-0400

"x^\frac mn=(x^m)^\frac 1n=\sqrt[n]{x^m}=(\sqrt[n]{x})^m."